Grade 12Physical Chemistry

An element(atomicmass=60 u) having face centered cubic unit cell has a density of 6.23g cm–³.what is the edge length of the unit cell?

Aomangyang
8 Years agoGrade 12

1 Answer

asit kumar
8 Years ago
d(density)=Z*M/(a)3*Na. given that:z=4(no of atoms per unit cell in fcc). M=60. d=6.23g/cm3,Na=6*10to the power23. putting values in the formula we will get the value of a(edge length).6.23=4*60/(a)3*6*10to the power23. we get (a)3=64*10 to the power-24 . so a=4*10to the power-8 ie4 angstrom.